OpenStudy (anonymous):

How do I simplify this to x^1 or just x?

OpenStudy (anonymous):

I know the answer, I just need to understand the steps

OpenStudy (anonymous):



OpenStudy (anonymous):

So yes your answer is x^1 = x.

OpenStudy (usukidoll):

also keep in mind that higher level math classes won't allow to write x^1... so it's best to know that x is just x to the first power .

OpenStudy (anonymous):

Yeah you can think of it this way, x^2 = x*x, x^3 = x*x*x, etc. So x^1 = just x. Hope that helps!

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @UsukiDoll also keep in mind that higher level math classes won't allow to write x^1... so it's best to know that x is just x to the first power . \(\color{blue}{\text{End of Quote}}\) Any algebra class will not allow it because \(x^1\) simplifies to \(x\) in the same way that \(x+0\) simplifies to \(x\). If you don't simplify, you will lose points.

OpenStudy (usukidoll):

and I'm still stuck on my stuff... sigh maybe I'll just go on a wing and a prayer when the site comes back up.

OpenStudy (anonymous):

It's just, how does 1/(x^-1) = x?

OpenStudy (anonymous):

It tells us that any nonzero number raised to a negative power equals its reciprocal raised to the opposite positive power.

OpenStudy (anonymous):

I see. Thanks.

OpenStudy (anonymous):

Haru, you want to know why \[ \frac 1x = x^{-1} \]?

OpenStudy (anonymous):

Does \[ x^{-1}=\frac 1x \]Make sense to you? Or does it seem weird?

OpenStudy (anonymous):

It wasn't decided completely arbitrarily. It is based off the following pattern: \[ x^a\times x^b=x^{a+b} \]